OpenStudy (anonymous):

17. Toby the cat is stuck in a tree. Billy's dad placed a ladder 5 feet from the base of the tree and then leaned the ladder against the tree. The top of the ladder is right where Toby is sitting. If the ladder makes a 65° angle with the ground, how far up the tree is Toby the cat? A. 9.3 feet B. 10.7 feet C. 11.8 feet D. 13.6 feet can someone help?

OpenStudy (anonymous):



OpenStudy (ajprincess):

 tan65=y/5 y=5*tan65. See if u can solve it.